# Market-Bench: Evaluating Large Language Models on Introductory Quantitative Trading and Market Dynamics

### v1

@@ 5.2 Strategy Difficulty
  • Strategy 1: Average pass@3 =0.80=0.80, average pass@1 =0.63=0.63, and average mean MAE $1.51\times 10^{3}$. Excluding Qwen3 Max, the average mean MAE drops to $1.43\times 10^{3}$.
  • Strategy 2: Average pass@3 =0.67=0.67, average pass@1 =0.47=0.47, and average mean MAE $5.21\times 10^{7}$, dominated by the extreme outlier from Qwen3 Max. Excluding Qwen3 Max, the average mean MAE falls to $2.61\times 10^{3}$.
  • Strategy 3: Average pass@3 =0.65=0.65, average pass@1 =0.52=0.52, and average mean MAE $5.57\times 10^{4}$. Excluding Qwen3 Max, the average mean MAE drops to $2.52\times 10^{4}$.
@@ 5.3 Overall Model Comparison
Table 4 presents overall metrics averaged across all three strategies, again sorted from best to worst by pass@3, pass@1, mean MAE, and average attempts.
Model pass@3 pass@1 Mean MAE Avg. attempts
gemini-3-pro-preview 1.00 0.93 437.51 1.07
gpt-5.1-codex-max 1.00 0.93 3,102.74 1.07
claude-sonnet-4.5 1.00 0.93 5,459.68 1.07
qwen3-max 1.00 0.80 190,973,360.19 1.20
mistral-large-2512 0.93 0.73 46,134.31 1.47
deepseek-v3.2 0.87 0.67 4,113.84 1.27
grok-4 0.87 0.47 705.16 1.67
llama-4-maverick 0.53 0.40 9,252.77 1.67
claude-opus-4.5 0.47 0.27 2,765.33 1.61
llama-3.1-nemotron-ultra 0.33 0.20 11,404.92 1.83
command-a 0.33 0.07 9,817.97 2.50
nova-premier-v1 0.13 0.07 688.57 1.50
@@
  • Reliability vs. accuracy. Qwen3 Max achieves perfect overall pass@3 ($1.00$) and high pass@1 ($0.80$), but its overall mean MAE is enormous ($\approx 1.91\times 10^{8}$) due to extreme divergence on Strategy 2. In contrast, Gemini 3 Pro combines perfect pass@3 with a very low mean MAE of 437.51, making it the best overall trade-off between reliability and numerical fidelity.
  • Inconsistency. Cohere Command-A, Amazon Nova Premier, and Nvidia Nemotron frequently fail to produce executable backtests on the harder strategies, leading to low overall pass@3 (0.33, 0.13, and 0.33, respectively) even when their MAE on the few solved rounds is not uniformly bad. Furthermore, almost every model showed large variations within the rounds themselves.
@@ 6 Discussion and Failure Modes
  • Miscomputing spreads or z-scores in Strategy 2 (e.g., mixing up hedge ratios between COKE and PEPSI or using inconsistent rolling windows),
  • Mishandling the options delta stream or hedge timing in Strategy 3, or interpreting delta signs incorrectly,
@@
Models that are consistent but logically unsound tend to show high pass@3 with large mean MAE, as seen with Qwen3 Max, Mistral-Large-2512, and, to a smaller extent, DeepSeek V3.2.

### v2

@@ 5.2 Strategy Difficulty
  • Strategy 1: Average Executable Passes =4.08=4.08 and average mean MAE $9.42\times 10^{2}$. Excluding Qwen3 Max, the average mean MAE drops to $8.53\times 10^{2}$.
  • Strategy 2: Average Executable Passes =3.46=3.46 and average mean MAE $3.15\times 10^{7}$, dominated by the extreme outlier from Qwen3 Max. Excluding Qwen3 Max, the average mean MAE falls to $4.50\times 10^{2}$.
  • Strategy 3: Average Executable Passes =3.38=3.38 and average mean MAE $4.20\times 10^{4}$. Excluding Qwen3 Max, the average mean MAE drops to $1.81\times 10^{4}$.
@@ 5.3 Overall Model Comparison
Table 4 presents overall metrics averaged across all three strategies, again sorted from best to worst by mean MAE.
Model Mean MAE Best Run MAE Executable Passes
grok-4 443.24 7.22 13
gpt-5.2 969.39 48.10 15
gemini-3-pro-preview 1,744.27 14.83 15
gpt-5.1-codex-max 4,242.93 0.002 15
deepseek-v3.2 4,575.64 7.26 13
claude-sonnet-4.5 5,126.76 16.35 15
claude-opus-4.5 6,039.62 7.18 7
command-a 6,562.11 133.10 5
nova-premier-v1 7,740.26 133.10 2
llama-3.1-nemotron-ultra 9,674.21 111.63 5
llama-4-maverick 10,202.18 131.92 8
mistral-large-2512 30,605.53 23.01 14
qwen3-max 159,144,490.12 16.39 15
@@
  • Reliability vs. accuracy. Qwen3 Max achieves perfect overall Executable Passes ($15$), but its overall mean MAE is enormous ($\approx 1.59\times 10^{8}$) due to extreme divergence on Strategy 2. In contrast, Grok 4 achieves the lowest overall mean MAE (443.24) despite not solving all rounds, and GPT-5.2 combines perfect executability with low mean MAE of 969.39.
  • Inconsistency. Cohere Command-A, Amazon Nova Premier, and Nvidia Nemotron frequently fail to produce executable backtests on the harder strategies, leading to low overall Executable Passes even when their MAE on the few solved rounds is not uniformly bad. Furthermore, almost every model showed large variations within the rounds themselves.
@@ 6 Discussion and Failure Modes
  • Miscomputing spreads or z-scores in Strategy 2 (for example, wrong hedge ratios between COKE and PEPSI or using wrong rolling windows),
  • Processing the options delta stream or hedge timing in Strategy 3 incorrectly, or interpreting delta signs incorrectly,
@@
Models that are consistent but logically unsound tend to show high Executable Passes with large mean MAE, as seen with Qwen3 Max, Mistral-Large-2512, and, to a smaller extent, DeepSeek V3.2.